OpenStudy (anonymous):

Sharmi and Reese each deposit money into a savings account. Sharmi deposits $300 less than Reese. After the same amount of time, Reese has earned $180 in interest at a rate of 3% per year, and Sharmi has earned $270 in interest at a rate of 6% per year. How much money did Reese and Sharmi initially deposit into their savings accounts?

OpenStudy (anonymous):

Here are the answers: Reese initially deposited $324, Sharmi initially deposited $24. Reese initially deposited $900, Sharmi initially deposited $600. Reese initially deposited $1,200, Sharmi initially deposited $900. Reese initially deposited $1,620, Sharmi initially deposited $1,300.

OpenStudy (mertsj):

x= amount Reese deposited x-300 = amount Sharmi deposited .03x = interest Reese earned .06(x-300)= interest Sharmi earned .03x+.06(x-300)=180+270